calculate the density of a solid which weighs 15N in air and 12N in water
Answer #1
what force causes every thing on earth to have wieght
answered by: John Ackah
Answer #2
Buoyancy = 3 N = rho of water * g * V
3 N = 1000 kg/m^3 * 9.81 m.s^2 * V
V = 2.06 *10^-4 m^3

mass = m = 15 N/ 9.81 = 1.53 kg
so
density = (1.53/2.06)*10^4 = 7423 kg/m^3
answered by: inny